OpenStudy (anonymous):

how to graph a function? For example, how do I graph the function d=3t. the problem says: You hike at 3 miles per hours. Your distance can be modeled by the function d=3t where t is time(hours). You plan on hiking for 10 hours. I also need to identify its domain and range.

OpenStudy (shamim):

here time t will b domain and distance d will b range

OpenStudy (shamim):

because time t is independent parameter and distance d is a dependent parameter

OpenStudy (anonymous):

Can I ask you a question on domain and range?

OpenStudy (shamim):

we can plot time t along x axis and distance d along y axis then u will get ur required graph

OpenStudy (shamim):

when time t=0 then distance d=0

OpenStudy (shamim):

t=1 hr d=3*1=3mile

OpenStudy (shamim):

t=2 hr d=3*2=6miles

OpenStudy (shamim):

t=3 hr d=3*3=9 miles

OpenStudy (shamim):

upto t= 10 hrs

OpenStudy (shamim):

t=10 hr d=3*10=30 miles

OpenStudy (shamim):

now u hv to plot all those value of both t and d.

OpenStudy (shamim):

do u need any more explanation abt ur graph plz ask me